40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Thu May 30 15:00:28 2013, Koji, Optics, General, QPD alignment 
    Reply  Fri May 31 05:46:54 2013, Koji, Optics, General, QPD alignment QPD1.JPGQPD2.JPGQPD_calib.nb.zip
Message ID: 132     Entry time: Thu May 30 15:00:28 2013     Reply to this: 133
Author: Koji 
Type: Optics 
Category: General 
Subject: QPD alignment 

The QPD alignment was adjusted using the aligned beam to the cavity and the 4ch transimpedance amplifier.

As I have a test cable for the QPD, I attached a DB9 connector on it so that I can use the QPD transimpedance
amplifier to read the photocurrent. The transimpedance of the circuit is 1kV/A.
As this board (D1001974) does not have X/Y/SUM output, I quickly made the summing circuit on a universal
board I took from Japan a while ago.

The spot on the QPD1 (shorter arm side) seems too low by ~0.64mm. It seems that the QPD is linearly responding
to the input misalignment, so there is no optical or electrical problem.

As I wonder how much I can improve the situation by replacing the diodes, I swapped the diodes between QPD1 and QPD2.
Now QPD1 and QPD2 have the diode #43 and #38, respectively. It improved the situation a llitle (about 60um).
But the beam is still 0.58mm too low. 95% of the power is on the upper two elements.
Of course this is at the edge of the linear range.
I confirmed we still can observe the cavity is fringing even with the beam is aligned on this QPD. So this may be
sufficient for the initial alignment.

The QPD2 was in a better situation. The spot is about 100um too low but this is still well with in the linear range.

The incident powers on the diodes were also measured. The estimated responsivities and Q.E.s are listed below.
The reflection from the diode is adjusted to hit the beam dump properly.

Here are the raw numbers


QPD#            QPD1       QPD2
Diode#          #43        #38

-------------------------------------
Power Incident  118.8 uW   115.7uW
Sum Out          78.8 mV   84.6 mV
Vertical Out     69.1 mV   11.9 mV
Horizontal Out    9.8 mV   -1.6 mV
SEG1             -1.90 mV -17.8 mV
SEG2             -2.18 mV -17.5 mV
SEG3            -32.0 mV  -25.3 mV
SEG4            -42.0 mV  -23.8 mV
-------------------------------------
Responsivity[A/W] 0.66      0.73
Q.E.              0.77      0.85
-------------------------------------

Arrangement of the segments
View from the beam
/ 2 | 1 X
|---+---|
\ 3 | 4 /

ELOG V3.1.3-